Problem 4.5

A bicycle tire company performed a web-based study of a popular tire retail price over time. The study indicated that price is set at $19.00 per tire, it was expected to increase to $26.00 over the next 5 years.

a. Determine the annual rate of inflation over 5 years to increase the price from $19.00 to $26.00. (10 points)

b. Determine the market interest rate that must be used in economic equivalence computations if inflation is considered and a 6% per year real interest rate is expected by this company. (10 points)
Answer #1

a.

using formula F = P*(1+i)^t

Let inflation rate be i%, then

26 = 19*(1+i)^5

(1+i)^5 = 26 / 19 = 1.368421

1+i = 1.368421^(1/5) = 1.064741

i = 0.064741 ~ 6.47%

b.

Market interest rate = 0.06 + 0.064741 + 0.06*0.064741 = 0.128625 ~ 12.86%